# Collision-Free Trajectory Optimization in Cluttered Environments with Sums-of-Squares Programming

### v1

@@ VI CONCLUSION
This paper introduces a novel framework for optimizing collision-free trajectories to navigate robots safely in cluttered environments. Our approach decomposes the free space into overlapping polytopic regions and formulates an SOS optimization problem to determine the minimum scaling factor for each region, such that the robot with specific geometry is contained within the scaled free region. We enforce the geometry-aware safety constraints by restricting that the scaling factor to be less than 1 along the entire trajectory. Importantly, the SOS programming is transformed into a convex SDP, which enables efficient computation of the solution and gradient information that are then integrated with AL-iLQR to generate effective and flexible collision-free trajectories. We extensively evaluate our proposed algorithm in both simulations and real-world experiments, and the results demonstrate that our method is capable of generating effective and flexible obstacle avoidance maneuvers when navigating the robot through challenging environments.

### v2

@@ VI CONCLUSION
This paper introduces a novel framework for optimizing collision-free trajectories to navigate robots safely in cluttered environments. Our approach decomposes the free space into overlapping polytopic regions and formulates an SOS optimization problem to determine the minimum scaling factor for each region, such that the robot with specific geometry is contained within the scaled free region. We enforce the geometry-aware safety constraints by restricting the scaling factor to be less than 1 along the entire trajectory. Gradient information of the SOS optimization problem is then integrated with AL-iLQR to generate collision-free trajectories. We extensively evaluate our proposed algorithm in both simulations and real-world experiments, and the results demonstrate that our method is capable of generating effective and flexible obstacle avoidance maneuvers when navigating the robot through challenging environments.